OpenStudy (anonymous):

Suppose z=x^2*sin(y), x=2s^2−4t^2, y=4st. A. Use the chain rule to find ∂z/∂s and ∂z/∂t as functions of x, y, s and t.

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

dz/ds = dz/dx*dx/ds + dz/dy*dy/ds

OpenStudy (anonymous):

dz/ds = 2x*sin(y)*(4s) + x^2*cos(y)*4t

OpenStudy (anonymous):

dz/dt = dz/dx*dx/dt + dz/dy*dy/dt

OpenStudy (anonymous):

dz/dt = -2x*sin(y)*8t + x^2*cos(y)*4s

OpenStudy (anonymous):

B. Find the numerical values of ∂z∂s and ∂z∂t when (s,t)=(−5,−1). ∂z∂s(−5,−1)= ∂z∂t(−5,−1)=

OpenStudy (anonymous):

substitute 2s^2 - 4t^2 for x, substitute 4st for y

OpenStudy (anonymous):

dz/ds in terms of s and t = 2*(2s^2 - 4t^2)*sin(4st)*(4s) + (2s^2 - 4t^2)^2*cos(4st)*(4s)

OpenStudy (anonymous):

Plug in -5 and -1 for t and s Respectively

OpenStudy (anonymous):

dz/ds (-5,-1) = -1840sin(20) - 8464(cos(20))

OpenStudy (anonymous):

dz/dt (-5, -1) = 8*92*sin(20) - 20*(46^2)*cos(20)

OpenStudy (anonymous):

dz/dt (-5, -1) = 736sin(20) - 42320cos(20)